Lori and Jason are gleaning (or collecting) avocados from a grove. Jason can glean 350 avocados per hour. Lori can glean 400 avocados per hour. If they start at the same time how long will it take both of them to collect a total of 10,500 avocados. Thank you!!
Answer by stanbon(75887)

Jason can glean 350 avocados per hour. Lori can glean 400 avocados per hour. If they start at the same time how long will it take both of them to collect a total of 10,500 avocados.
----
Each hour they are gleaning (350+400) = 750 avocados.
----
time to glean 10,500 = 10500/750 = 14 hrs.
==============================================
Cheers,
Stan H.
============